# Geometric Algebra Planes: Convex Implicit Neural Volumes

### v1

@@ 4.1 Equivalence to Matrix Completion in 2D
We can interpret the matrix completion model of Equation 12 as a sum of $h$ different low-rank approximations, where the matrices within each of the $h$ groups are constrained to share the same singular vectors $U_{j},V_{j}$. The binary masks $B_{i}$ effectively allow each of these $h$ low-rank approximations to attend to (or complete) a different part of the matrix $M$ before being linearly combined through the weights (singular values) $W_{i,j}$. The upper limit of the rank of this matrix approximation is thus $hk$, fundamentally larger than the rank constraint faced by models with a linear decoder.
@@
where $W\in\mathbb{R}{h\times k}$ is the weight matrix for the MLP decoder’s hidden layer (with width $h$) and $\alpha\in\mathbb{R}^{h}$ is the weight vector of the MLP decoder’s output layer.
@@
The upper limit of the rank of this matrix approximation is $hk$, the same as with a convex MLP decoder.
We summarize the maximum attainable ranks of different 2D models in Table 1 (see Section A.2.3 and Section A.2.4 for matrix representations of $\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ and $\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$ with convex and nonconvex MLP decoders). Experimental validation of these theoretical results on the task of 2D image fitting (compression) is provided in Figure 2, with a comparison of interpolation schemes in Figure 8 in the appendix. As expected, we find that a linear decoder model with multiplication dramatically outperforms its additive counterpart, which does not improve with increasing model size. We also find that 2D GA-Planes models with MLP decoders can match or exceed the compression performance of the optimal low-rank representation found by singular value decomposition (SVD), especially when using a nonconvex MLP. This is a testament to the capacity of an MLP decoder to increase representation rank using fewer parameters than a traditional low-rank decomposition, as well as to the resolution compressibility of natural images. These experimental results also provide complementary information the one-sided bounds on representation rank and fitting error in our theoretical analysis. For example, with an MLP decoder, the theoretical maximum rank of a $\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ model exceeds the theoretical maximum rank of a $\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})}$ model. However, this gap is not realized in practice because the additive feature model fails to achieve its theoretical maximum rank.
Model Linear decoder MLP decoder convex MLP decoder
$\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ $2$ $2hk$ $2hk$
$\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$ $2$ $2hk$ $2hk$
$\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})}$ $k$ $hk$ $hk$
Table 1: Maximum attainable ranks of different 2D GA-Planes models, using only line features. Here $h$ denotes hidden decoder dimension and $k$ denotes feature dimension. Replacing a linear decoder with a convex or nonconvex MLP can dramatically increase the rank of the representation.
@@ 4.2 Lower Bounds
Based on the matrix completion theorems and their summary in Table 1, we present lower bounds on the feature dimension (rank) $k$ necessary to attain zero representation error for different models. To achieve zero error with a 2D GA-Planes model $\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})}$, $k$ must satisfy
$k>\frac{\min(m,n)}{h},$ (14)
where $m,n$ are the dimensions of the target image (assumed to have full rank) and $h$ is the hidden dimension of the convex or nonconvex MLP decoder ($h=1$ for linear decoder). Similarly, when the model is $\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ or $\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$, and the decoder is a convex or nonconvex MLP, $k$ must satisfy
$k>\frac{min(m,n)}{2h}.$ (15)
If this model uses a linear decoder, its representation has rank $2$ no matter the value of $k$. Increasing the feature dimension $k$ in this setting does not improve the rank capacity of the representation.
In some cases, we do not require our representation to achieve zero error. If we instead wish to fit up to the $j$-th singular value of $M$ (sorted in decreasing order), $k$ must satisfy
$k>\frac{j}{h},$ (16)
for $\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})}$,
$k>\frac{j}{2h},$ (17)
for $\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ or $\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$. Again, these bounds hold for all model combinations in Table 1 except for the linear decoder operating on added or concatenated features; in this case there is no benefit to increasing $k$.
Finally, we express these results in terms of lower bounds on the Frobenius norm errors of each 2D GA-Planes model. We denote the optimal fitting error of the linear and MLP decoder models by $E_{linear}(\text{D}{(f(q))})$ and $E_{MLP}(\text{D}{(f(q))})$ for different feature combinations $f(q)$. For models with a linear decoder,
@@
$E_{MLP}(\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})})$ $\geq\sigma_{2hk}(M)$ (21)
$E_{MLP}(\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})})$ $\geq\sigma_{hk}(M)$ (22)
$E_{MLP}(\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2}+\mathbf{e}_{12})})$ $\geq\sigma_{hk}(M-\varphi(L^{*})).$ (23)
@@ 5.1 Radiance Field Modeling
Our experiments for the 3D segmentation task use the opacity masks from the NeRF-Blender lego scene (Mildenhall et al., 2020). The task is to “lift” 2D segmentation masks to 3D, rather than generating 3D segmentations directly from raw photographs. We compare the GA-Planes architecture with concatenation of features (described by eq. 4) and the simpler Tri-Plane representation proposed in Chan et al. (2022), in which the three plane features are added together and decoded without use of line or volume features ($\text{D}(\mathbf{e}_{12}+\mathbf{e}_{13}+\mathbf{e}_{23})$). We train three versions of each model: convex, semiconvex, and nonconvex, to validate that the GA-Planes architecture in eq. 4 is more robust than the Tri-Plane model across these different formulations.
@@
Tri-Planes (with $+$) 0.681 0.868 0.863
@@
We find that the GA-Planes architecture outperforms the Tri-Planes architecture when using identical feature resolution and dimensions, regardless of whether training is convex, semiconvex, or nonconvex. Further, we see that GA-Planes retains similar performance regardless of optimization strategy, whereas the Tri-Plane model learns poorly via convex optimization.
@@
Our second set of 3D segmentation experiments leverages direct 3D supervision via Space Carving (Kutulakos & Seitz, 1999). Space Carving supervision operates on the principle that if any ray passing through a given 3D coordinate is transparent, the density at that 3D coordinate must be zero. This method recovers the visual hull of the object by “carving out” empty space around it. Our results with 3D supervision, in Table 3, parallel those with 2D supervision: GA-Planes performs well regardless of whether it is optimized in a convex, semiconvex, or nonconvex formulation, whereas the Tri-Planes model performs decently under nonconvex optimization but much worse with convex or semiconvex optimization.
@@
Tri-Planes (with $+$) 0.642 0.636 0.941
@@ 5.3 Video Segmentation
Our results are summarized in Figure 7 and Table 4. Similar to the volume segmentation setting, here we again find that GA-Planes performs well across convex, semiconvex, and nonconvex training, though its performance is slightly reduced under fully convex training, perhaps because the convex model size is slightly reduced due to fusing the decoder parameters into the feature grids. In contrast, the simpler Tri-Plane models perform poorly on this task regardless of training strategy: they fail to learn the temporal sequence of the video, producing masks that focus only on the skateboarder’s less-mobile core. In this experiment, we also compare nonconvex models of each type (Tri-Plane and full GA-Planes) in which the features are combined in a linear way (by concatenation or addition) versus a nonlinear way (by multiplication). Although in general (from our theory) we expect multiplication of features to be more expressive than either addition or concatenation, we do not find a meaningful difference in this experiment. This similar performance between addition/concatenation and multiplication of features is in line with results reported in a similar ablation study in K-Planes (Fridovich-Keil et al., 2023), in which the benefits of feature multiplication were only evident when using a linear decoder rather than a nonlinear MLP decoder as is used here.
@@ Acknowledgments
This work was supported in part by the NSF Mathematical Sciences Postdoctoral Research Fellowship under award number 2303178, in part by the National Science Foundation (NSF) under Grant DMS-2134248, in part by the NSF CAREER Award under Grant CCF-2236829, in part by the U.S. Army Research Office Early Career Award under Grant W911NF-21-1-0242, and in part by the Office of Naval Research under Grant N00014-24-1-2164.
@@ A.2 Proof of Theorems
In order to represent a matrix $M\in\mathbb{R}^{m\times n}$ with an implicit model, we compute $\text{D}{(f(q))}$ for $q=(k,l),\,\ \forall k\in\{1,\dots,m\},\,\ \forall l\in\{1,\dots,n\}$. Considering line feature grids with resolutions matching $m$, $n$; the features will become $\mathbf{e}_{1}={(\mathbf{g}_{1})}_{k}$, $\mathbf{e}_{2}={(\mathbf{g}_{2})}_{l}$ for $q=(k,l)$ otherwise they will be $\mathbf{e}_{1}={\varphi(\mathbf{g}_{1})}_{k}$, $\mathbf{e}_{2}={\varphi(\mathbf{g}_{2})}_{l}$ where $\varphi(\mathbf{g}_{1})$, $\varphi(\mathbf{g}_{2})$ now have resolutions $m$, $n$ after interpolation through $\varphi$. Here $\varphi$ can be any interpolation scheme with linear weighting of inputs, e.g. nearest neighbor, (bi)linear, (bi)cubic, spline, Gaussian, sinc, etc. For the simplicity of notation, we omit $\varphi$ in line feature grids, and only apply it to the plane feature grid $\mathbf{g}_{12}$, which has lower resolution by design. The proofs consider $\mathbf{g}_{1},\mathbf{g}_{2}\in\mathbb{R}^{r_{1}\times d_{1}}$ and $\mathbf{g}_{12}\in\mathbb{R}^{r_{2}\times r_{2}\times d_{1}}$ (equal feature dimensions, different resolutions), resulting in the matrix representation $\hat{M}\in\mathbb{R}^{r_{1}\times r_{1}}$ (the case where $m=n=r_{1}$). We derive expressions for $\hat{M}$ implied by different GA-Planes variations in the parts that follow. The coordinate-wise optimization objective (in the case of a directly supervised mean-square-error loss) corresponds to minimizing the Frobenius norm of the ground truth matrix $M$ and its approximation $\hat{M}$.
@@ A.2.3 Proof of theorem 3
It can be observed from eq. 45 that the additive-feature model with a convex MLP decoder enjoys the increased maximum rank capacity of $2hd_{1}$, though in practice this maximum capacity may not always be achieved (see Figure 2).
@@ A.2.4 Proof of theorem 4
For a standard 2-layer ReLU MLP with hidden size $h$, denote the trainable first and second layer weights as $W\in\mathbb{R}^{h\times d_{1}}$, $\alpha\in\mathbb{R}^{h\times 1}$. We will handle three different cases for merging the interpolated features: multiplication ($\circ$), addition ($+$), and concatenation ($odot$).
@@ A.4 Benefits of Convexity
In most of our experiments, all models (convex, semiconvex, and nonconvex) are large enough that they are able to optimize well. However, we highlight a benefit of our convex and semiconvex models that they enjoy more stable optimization even with very small model sizes. In Figure 9 we compare training curves for very small models for our video fitting task (hidden dimension 4 in the decoder MLP, and feature dimensions $[d_{1},d_{2},d_{3}]=[4,4,2]$ and resolutions $[r_{1},r_{2},r_{3}]=[32,32,16]$ for line, plane, and volume features, respectively). We repeat optimization with 3 different random seeds used to initialize the optimizable parameters (gating weights for the convex and semiconvex models are fixed). While the convex and semiconvex models enjoy stable training curves across random seeds, we find that the nonconvex model experiences much more volatile training behavior (completely failing to optimize with one of the 3 random seeds).
Figure 9: Training curves for a very small GA-Planes model on our video fitting task, across 3 random seeds. We find that the favorable optimization landscape of our convex and semiconvex models enables reliable training across seeds, whereas the nonconvex model fails to train with one of the seeds.

### v2

@@ 4.1 Equivalence to Matrix Completion in 2D
We can interpret the matrix completion model of Equation 12 as a sum of $h$ different low-rank approximations, where the matrices within each of the $h$ groups are constrained to share the same singular vectors $U_{j},V_{j}$. The binary masks $B_{i}$ effectively allow each of these $h$ low-rank approximations to attend to (or complete) a different part of the matrix $M$ before being linearly combined through the weights (singular values) $W_{i,j}$. The upper limit of the rank of this matrix approximation is thus $\min(n,m)$, because the mask matrices can arbitrarily increase the rank beyond the constraint faced by models with a linear decoder. Note that if the feature grids $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$ have spatial resolution $r_{1}$ less than $\min(n,m)$, the maximum rank will be $r_{1}$.
@@
where $W\in\mathbb{R}^{h\times k}$ is the weight matrix for the MLP decoder’s hidden layer (with width $h$) and $\alpha\in\mathbb{R}^{h}$ is the weight vector of the MLP decoder’s output layer.
@@
The upper limit of the rank of this matrix approximation is $\min(n,m)$, the same as with a convex MLP decoder. Again, note that if the feature grids $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$ have spatial resolution $r_{1}$ less than $\min(n,m)$, the maximum rank will be $r_{1}$.
We summarize the maximum attainable ranks of different 2D models in Table 1 (see Section A.2.3 and Section A.2.4 for matrix representations of $\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ and $\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$ with convex and nonconvex MLP decoders). Experimental validation of these theoretical results on the task of 2D image fitting (compression) is provided in Figure 2, with a comparison of interpolation schemes in Figure 8 in the appendix. As expected, we find that a linear decoder model with multiplication dramatically outperforms its additive counterpart, which does not improve with increasing model size. We also find that 2D GA-Planes models with MLP decoders can match or exceed the compression performance of the optimal low-rank representation found by singular value decomposition (SVD), especially when using a nonconvex MLP. This is a testament to the capacity of an MLP decoder to increase representation rank using fewer parameters than a traditional low-rank decomposition, as well as to the resolution compressibility of natural images. These experimental results also provide complementary information the one-sided bounds on representation rank and fitting error in our theoretical analysis.
Model Linear decoder convex MLP decoder MLP decoder
$\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})}$ $2$ $r_{1}$ $r_{1}$
$\text{D}{(\mathbf{e}_{1}\odot\mathbf{e}_{2})}$ $2$ $r_{1}$ $r_{1}$
$\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})}$ $k$ $r_{1}$ $r_{1}$
Table 1: Maximum attainable ranks of different 2D GA-Planes models, using only line features. Here $k$ is the feature dimension and $r_{1}$ is the spatial dimension of the features, which need never exceed $\min(m,n)$. Replacing a linear decoder with a convex or nonconvex MLP can dramatically increase the rank of the representation.
@@ 4.2 Lower Bounds
Based on the matrix completion theorems and their summary in Table 1, we present lower bounds on the Frobenius norm errors of each 2D GA-Planes model. We denote the optimal fitting error of the linear and MLP decoder models by $E_{linear}(\text{D}{(f(q))})$ and $E_{MLP}(\text{D}{(f(q))})$ for different feature combinations $f(q)$. For models with a linear decoder,
@@
$E_{MLP}(\text{D}{(\mathbf{e}_{1}+\mathbf{e}_{2})})$ $\geq\sigma_{r_{1}}(M)$ (17)
$E_{MLP}(\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2})})$ $\geq\sigma_{r_{1}}(M)$ (18)
$E_{MLP}(\text{D}{(\mathbf{e}_{1}\circ\mathbf{e}_{2}+\mathbf{e}_{12})})$ $\geq\sigma_{r_{1}}(M-\varphi(L^{*})).$ (19)
@@ 5.1 Radiance Field Modeling
Our experiments for the 3D segmentation task use the opacity masks from the NeRF-Blender lego scene (Mildenhall et al., 2020). The task is to “lift” 2D segmentation masks to 3D, rather than generating 3D segmentations directly from raw photographs. We compare the GA-Planes architecture with concatenation of features (described by eq. 4) and the simpler Tri-Plane representation proposed in Chan et al. (2022), in which the three plane features are added together and decoded without use of line or volume features ($\text{D}(\mathbf{e}_{12}+\mathbf{e}_{13}+\mathbf{e}_{23})$). We train three versions of each model: convex, semiconvex, and nonconvex, to validate that the GA-Planes architecture in eq. 4 is more robust than the Tri-Plane model across these different formulations. We also compare GA-Planes with multiplication of features, in eq. 5, to the Tri-Plane model using multiplication of plane features, like K-Planes ($\text{D}(\mathbf{e}_{12}\circ\mathbf{e}_{13}\circ\mathbf{e}_{23})$). These only have nonconvex formulations.
@@
Tri-Planes (planes with $\circ$, like K-Planes) - - 0.877
Tri-Planes (planes with $+$, like (Chan et al., 2022)) 0.681 0.868 0.863
@@
We find that the GA-Planes architecture with concatenation outperforms the Tri-Planes architecture using addition when trained with equal total number of parameters (see Section A.7 for specific feature resolution and dimensions), regardless of whether training is convex, semiconvex, or nonconvex. Further, we see that GA-Planes retains similar performance regardless of optimization strategy, whereas the Tri-Plane model learns poorly via convex optimization. When the features are combined through multiplication, the models achieve the same IOU score.
@@
Our second set of 3D segmentation experiments leverages direct 3D supervision via Space Carving (Kutulakos & Seitz, 1999). Space Carving supervision operates on the principle that if any ray passing through a given 3D coordinate is transparent, the density at that 3D coordinate must be zero. This method recovers the visual hull of the object by “carving out” empty space around it. Our results with 3D supervision, in Table 3, parallel those with 2D supervision: GA-Planes performs well regardless of whether it is optimized in a convex, semiconvex, or nonconvex formulation, whereas the Tri-Plane model performs decently under nonconvex optimization but much worse with convex or semiconvex formulation.
@@
Tri-Planes (planes with $\circ$, like K-Planes) - - 0.881
Tri-Planes (planes with $+$, like (Chan et al., 2022)) 0.642 0.636 0.941
@@ 5.3 Video Segmentation
Our results are summarized in Figure 7 and Table 4. Similar to the volume segmentation setting, here we again find that GA-Planes performs well across convex, semiconvex, and nonconvex training, though its performance is slightly reduced under fully convex training, perhaps because the convex model size is slightly reduced due to fusing the decoder parameters into the feature grids. In contrast, the simpler Tri-Plane models perform poorly on this task regardless of training strategy: they fail to learn the temporal sequence of the video, producing masks that focus only on the skateboarder’s less-mobile core. In this experiment, we also compare nonconvex models of each type (Tri-Plane and full GA-Planes) in which the features are combined in a linear way (by concatenation or addition) versus a nonlinear way (by multiplication). The similar performance between addition/concatenation and multiplication of features observed here is in line with results reported in a similar ablation study in K-Planes (Fridovich-Keil et al., 2023), in which the benefits of feature multiplication were only evident when using a linear decoder rather than a nonlinear MLP decoder as is used here.
@@ Acknowledgments
We are grateful to Axel Levy for helpful discussions on space carving, and to Abhishek Shetty for helpful discussions on matrix completion. This work was supported in part by the NSF Mathematical Sciences Postdoctoral Research Fellowship under award number 2303178, in part by the National Science Foundation (NSF) under Grant DMS-2134248, in part by the NSF CAREER Award under Grant CCF-2236829, in part by the U.S. Army Research Office Early Career Award under Grant W911NF-21-1-0242, and in part by the Office of Naval Research under Grant N00014-24-1-2164.
@@ A.2 Proof of Theorems
In order to represent a matrix $M\in\mathbb{R}^{m\times n}$ with an implicit model, we compute $\text{D}{(f(q))}$ for $q=(k,l),\,\ \forall k\in\{1,\dots,m\},\,\ \forall l\in\{1,\dots,n\}$. Considering line feature grids with resolutions matching $m$, $n$; the features will become $\mathbf{e}_{1}={(\mathbf{g}_{1})}_{k}$, $\mathbf{e}_{2}={(\mathbf{g}_{2})}_{l}$ for $q=(k,l)$ otherwise they will be $\mathbf{e}_{1}={\varphi(\mathbf{g}_{1})}_{k}$, $\mathbf{e}_{2}={\varphi(\mathbf{g}_{2})}_{l}$ where $\varphi(\mathbf{g}_{1})$, $\varphi(\mathbf{g}_{2})$ now have resolutions $m$, $n$ after interpolation through $\varphi$. Here $\varphi$ can be any interpolation scheme with linear weighting of inputs, e.g. nearest neighbor, (bi)linear, (bi)cubic, spline, Gaussian, sinc, etc. For the simplicity of notation, we omit $\varphi$ in line feature grids, and only apply it to the plane feature grid $\mathbf{g}_{12}$, which has lower resolution by design. The proofs consider $\mathbf{g}_{1},\mathbf{g}_{2}\in\mathbb{R}^{r_{1}\times d_{1}}$ and $\mathbf{g}_{12}\in\mathbb{R}^{r_{2}\times r_{2}\times d_{1}}$ (equal feature dimensions, different resolutions), resulting in the matrix representation $\hat{M}\in\mathbb{R}^{r_{1}\times r_{1}}$ (the case where $m=n=r_{1}$). Note that if the interpolation is done by a method other than nearest neighbor, this may allow a (convex or nonconvex) MLP decoder to increase the rank beyond $r_{1}$. We derive expressions for $\hat{M}$ implied by different GA-Planes variations in the parts that follow. The coordinate-wise optimization objective (in the case of a directly supervised mean-square-error loss) corresponds to minimizing the Frobenius norm of the ground truth matrix $M$ and its approximation $\hat{M}$.
@@ A.2.3 Proof of theorem 3
In all these representations, a low-rank matrix is multiplied elementwise with a binary mask $B_{i}$, which makes the maximum attainable rank $r_{1}$. Thus, with a convex MLP decoder, rank of $\hat{M}$ is limited by the resolution of the feature grids.
@@ A.2.4 Proof of theorem 4
For a standard 2-layer ReLU MLP with hidden size $h$, denote the trainable first and second layer weights as $W\in\mathbb{R}^{h\times d_{1}}$, $\alpha\in\mathbb{R}^{h\times 1}$. We will handle three different cases for merging the interpolated features: multiplication ($\circ$), addition ($+$), and concatenation ($\odot$).
@@
By a similar argument to Section A.2.3, the maximum attainable rank of all three representations derived here is limited by $r_{1}$.
@@ A.4 Benefits of Convexity
In most of our experiments, all models (convex, semiconvex, and nonconvex) are large enough that they are able to optimize well. However, we highlight a benefit of our convex and semiconvex models that they enjoy more stable optimization even with very small model sizes. In Figure 9 we compare test intersection-over-union (IoU) curves for very small models for our video fitting task (hidden dimension 4 in the decoder MLP, and feature dimensions $[d_{1},d_{2},d_{3}]=[4,4,2]$ and resolutions $[r_{1},r_{2},r_{3}]=[32,32,16]$ for line, plane, and volume features, respectively). We repeat optimization with 3 different random seeds used to initialize the optimizable parameters (gating weights for the convex and semiconvex models are fixed). While the convex and semiconvex models enjoy stable training curves across random seeds, we find that the nonconvex model experiences much more volatile training behavior (completely failing to optimize with one of the 3 random seeds).
Figure 9: Test performance throughout training on our video fitting task, for a very small GA-Planes model across 3 random seeds. We find that the favorable optimization landscape of our convex and semiconvex models enables reliable training across seeds, whereas the nonconvex model fails to fit any test frame with one of the seeds.